# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ 8.3. Totally-nested tetra-separations from good bags
Case $\pi=\{x,y\}$. We may assume that both $z_{1},z_{2}$ are vertices, by (6.4.1). Since the bag $G_{t(0)}$ is 2-connected, it contains two independent $x^{\prime}$$y^{\prime}$ paths. By Lemma 8.21, these paths contain two independent $\hat{x}$$\hat{y}$ paths which are also included in the bag $G_{t(0)}$. We find a third $\hat{x}$$\hat{y}$ path $G$ that avoids both the bag $G_{t(0)}$ and the centre, by the (One-Path Lemma). (8.10). So we are done by (6.4.3).
Case $\pi=\{z_{1},z_{2}\}$. We find three independent $z^{\prime}_{1}$$z^{\prime}_{2}$ paths in $G-x-y$ by Lemma 8.19. These contain three independent $\hat{z}_{1}$$\hat{z}_{2}$ paths by Lemma 8.21. So we are done by (6.4.3).
Case $\pi=\{x,z_{1}\}$. We find an $x^{\prime}$$z^{\prime}_{1}$ path with interior in $\mathring{G}_{t(0)}$, by the (Two-Fan Lemma). (8.11). This path contains a $\hat{x}$$\hat{z}_{1}$ path $P_{1}$, by Lemma 8.21. Next, we find two independent $x^{W}$$z^{\prime}_{1}$ paths $P_{2}$ and $P_{3}$ in $G$ that internally avoid $G_{t(0)}$, by Lemma 8.20. Both paths $P_{2}$ and $P_{3}$ are $\hat{x}$$\hat{z}_{1}$ paths by Lemma 8.16. Hence the $P_{i}$ witness (6.4.3).
@@
We claim that $z^{\prime}_{1}=z_{1}$. On the one hand, $z^{\prime}_{1}$ has $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma 8.16, so $z^{\prime}_{1}$ remains in the separator when taking the right-shift $(U^{\prime\prime},W^{\prime\prime})$ of $(U^{\prime},W^{\prime})$. However, $y^{\prime}$ becomes the edge $y$ of the $K_{2}$-bag $G_{t(1)}$ in the separator $S(U^{\prime\prime},W^{\prime\prime})$ by 8.15. Hence $z^{\prime}_{1}$ has two neighbours $y^{\prime}$ and $v$ in $U^{\prime\prime}\smallsetminus W^{\prime\prime}$. Thus $z^{\prime}_{1}$ also remains in the separator when taking the left-shift $(U,W)$ of $(U^{\prime\prime},W^{\prime\prime})$. In particular, $z^{\prime}_{1}=z_{1}$.
@@
Suppose otherwise that $G_{t(0)}$ is 2-connected and has at least four vertices. We will find five independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G$ for (6.7.2). There exist two disjoint $\{z^{\prime}_{1},z^{\prime}_{2}\}$$\mathring{G}_{t(0)}$ paths $P_{1},P_{2}$ by Lemma 8.8. Since $G_{t(0)}$ is 2-connected, we can extend the paths $P_{1},P_{2}$ to two disjoint $\{z^{\prime}_{1},z^{\prime}_{2}\}$$\{x^{\prime},y^{\prime}\}$ paths with interior in $\mathring{G}_{t(0)}$. We find three independent $\{x_{1}^{W},x_{2}^{W}\}$$\{z_{1}^{\prime},z_{2}^{\prime}\}$ paths in $G[W]$ by Lemma 8.20. The five paths are independent, and they contain five $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths by Lemma 8.21.
Case $\pi_{1}=\{x,z_{1}\}$ and $\pi_{2}=\{y,z_{2}\}$. Here it suffices to find five independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G$ for (6.7.2). There are two independent $x^{\prime}$$y^{\prime}$ paths in $G_{t(0)}$ by 2-connectivity. We find three independent $z^{\prime}_{1}$$z^{\prime}_{2}$ paths in $G[W]-\{x^{\prime},y^{\prime}\}$ by Lemma 8.19. The latter three paths are internally disjoint from the former two. By Lemma 8.21, the five paths contain five independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths.
@@
Assume $\pi=\{x,z_{1}\}$. The vertex $z^{\prime}_{1}$ in the centre has $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma 8.16. Hence $z^{W}_{1}=z^{\prime}_{1}$. Since the $\geqslant 3$ neighbours also lie in $W\smallsetminus U$ by 5.9, we further get that $z_{1}$ must be the edge in $\pi$. Since $z_{1}$ is an edge, $x$ is a vertex which has exactly one neighbour in $G[W]-z_{1}^{W}$. Hence $G_{t(-1)}$ is a $K_{2}$. But then $x$ is the edge of $G_{t(-1)}$ by 8.15, a contradiction.
Assume $\pi=\{x,y\}$. Say $x$ is a vertex and $y$ is an edge. Then $x=x^{\prime}$, so $G_{t(-1)}$ is 2-connected by 8.15. Hence $x$ has two neighbours $u,v$ in $G_{t(-1)}$. By the (Helpful Lemma). (8.6), neither $u$ nor $v$ is an endvertex of the edge $y$. Hence $\{x,y\}$ is not $W$-weird, a contradiction.
@@
Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$. Neither $\pi_{i}$ is $W$-weird by Lemma 8.25. In particular, $\{\pi_{1},\pi_{2}\}$ is 2-potter-linked.
To show that $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked, it suffices to find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$. If one of the $\pi_{i}$ is of the form $\{x,y\}$, then we find three such paths by Lemma 8.20 and Lemma 8.21. Otherwise $z_{1}\in\pi_{1}$ and $z_{2}\in\pi_{2}$, say. Then we find three desired paths by Lemma 8.19 and Lemma 8.21. ∎
@@
We may assume the (Block-bagel bag setting). (8.14) with $t=t(0)$. Then $(U_{t},W_{t})$ is a tetra-separation by Lemma 8.17. It suffices to show that $(U_{t},W_{t})$ is externally 5-connected, by 6.16. This follows from Lemma 8.22 (half-connected), Lemma 8.23 (3-linked), Lemma 8.24 (0-potter-linked) and Lemma 8.26 (1-potter-linked and 2-potter-linked). ∎
@@ 8.4. Splitting stars and torsos
The following two lemmata are supported by Figure 8.7.
@@
Assume the (Block-bagel setting). (8.4). Let $P$ and $Q$ be the two paths that are the components of $O-o(C)-o(D)$. Assume $o(A)\in P$ and $o(B)\in Q$. Let $c$ denote the end of $P$ that is adjacent or incident to $o(C)$ in $O$. Define $d$ similarly. Let $s$ be a node in the same component of $P-o(A)$ as $c$ such that $G_{s}=K_{2}$. Let $P^{\prime}:=dP\mathring{s}$ and $Q^{\prime}:=\mathring{s}Pc\cup Q$.
@@
Matching-condition: We have to show that, if $d$ in an edge, then it does not share an endvertex with $e$. The endvertices of $d$ lie in the corners for $AD$ and $BD$ since there do not exist dangling edges in the crossing-diagram of $(A,B)$ and $(C,D)$ by the (Crossing Lemma). (4.1). Both endvertices of $e$ lie in the corner for $AC$ by Lemma 8.2 (i). Hence $d$ and $e$ share no ends.
@@
If $d$ is a vertex, then we claim that every neighbour $u$ of $d$ in $A\smallsetminus B$ also also lies in $A^{\prime}\smallsetminus B^{\prime}$. Indeed, since $ud$ is not a dangling edge, $u$ is contained in the corner $AD$ or the $A$-link with respect to $(A,B),(C,D)$. By Lemma 8.2 (i), no endvertex of $e$ lies in the $A$-link with respect to $(A,B),(C,D)$. So, we have $u\in A^{\prime}\smallsetminus B^{\prime}$.
It remains to show that both vertices $z_{1}^{\prime}$ and $z_{2}^{\prime}$ in the centre have at least two neighbours in $A^{\prime}\smallsetminus B^{\prime}$. Let $s^{\prime}$ be the neighbour of $s$ on $P^{\prime}$. Observe that either $s^{\prime}$ or $ss^{\prime}$ is helpful and belongs to the corner $AC$. Moreover, by the (Helpful Lemma). (8.6) there is a helpful element that belongs to the $AD$-corner. Hence, there are at least two helpful elements that belong to $(ss^{\prime})\cup T_{AD}$, where $T_{AD}$ is the $AD$-block-path. By Lemmas 8.7 and 8.8, we are done.
@@
Assume the (Block-bagel setting). (8.4). Let $P$ and $Q$ be the two paths that are the components of $O-o(C)-o(D)$. Assume $o(A)\in P$ and $o(B)\in Q$. Let $c$ denote the end of $P$ that is adjacent or incident to $o(C)$ in $O$. Define $d$ similarly. Let $e$ be an edge of $O$ with 2-connected bags at both ends. Assume that $e$ stems from the $AC$-block-path. Let $P^{\prime}:=dP\mathring{e}$ and $Q^{\prime}:=\mathring{e}Pc\cup Q$. Then:
@@
If the $A$-link or the $C$-link contains a vertex, then this vertex is contained in a 2-connected bag $G_{t}$ with $t\in V(T^{\prime})$, by Lemma 8.2. Otherwise both the $A$-link and the $C$-link consist of edges. Then $G_{o(A)}$ and $G_{o(C)}$ are $K_{2}$-bags. Each of the two bags $G_{o(A)},G_{o(C)}$ has exactly one endvertex $v_{A},v_{C}$ in the corner for $AC$ by Lemma 8.2, respectively. If $T^{\prime}$ is empty, then $v_{A}=v_{C}$, and $A\cap C=\{v_{A}\}\cup Z^{\prime}$ has size three. ∎
@@
Note that the existence of $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ follows from $(A,B)$ and $(C,D)$ exhibiting properties (i)(iii). Let $T$ denote the $A^{\prime}C^{\prime}$-block-path, and let $T^{\prime}:=T-o(A^{\prime})-o(C^{\prime})$. By Lemma 8.32, $T^{\prime}$ is non-empty. Let $\mathcal{O}^{\prime}=\mathcal{O}$ be the block-bagel induced by $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$.
@@
[…]
@@
[…]
@@
Assume for contradiction that $\sigma(\mathcal{O})$ is interlaced by a totally-nested tetra-separation $(E,F)$ of $G$. By Lemma 8.33, there is a 2-connected bag $G_{t}$ of $\mathcal{O}$ with $E\subseteq U_{t}^{\prime}$, say. By Lemma 5.7, we have $(E,F)\leqslant(U_{t}^{\prime\prime},W_{t}^{\prime\prime})$ where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ is the right-shift of $(U^{\prime}_{t},W^{\prime}_{t})$.
Assume first that $G_{t}$ is good. Then $(U_{t},W_{t})\in\sigma(\mathcal{O})$ by definition of $\sigma(\mathcal{O})$. Since $(E,F)$ interlaces $\sigma(\mathcal{O})$, we have $(U_{t},W_{t})<(E,F)$ or $(U_{t},W_{t})<(F,E)$. We cannot have $(U_{t},W_{t})<(F,E)$ since otherwise $\emptyset\neq E\smallsetminus F\subseteq(W_{t}\smallsetminus U_{t})\cap(U_{t}^{\prime\prime}\smallsetminus W_{t}^{\prime\prime})$ contradicts that $U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}\subseteq U_{t}\smallsetminus W_{t}$ by 5.2. So $(U_{t},W_{t})<(E,F)\leqslant(U_{t}^{\prime\prime},W_{t}^{\prime\prime})$. Since $(U_{t},W_{t})$ is the left-shift of $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, some vertex in $S(E,F)$ must have $\leqslant 1$ neighbour in $E\smallsetminus F$, contradicting the degree-condition.
Otherwise, $G_{t}$ is bad. Since $G_{t}$ is 2-connected, it is a triangle, and since $G_{t}$ is bad, both neighbouring bags in $\mathcal{O}$ are 2-connected. Hence both adhesion-vertices of $G_{t}$ are contained in the separator of $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$. Therefore, the unique interior vertex of the triangle-bag $G_{t}$ is the only vertex of $U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}$. Since $(E,F)\leqslant(U_{t}^{\prime\prime},W_{t}^{\prime\prime})$, we have $E\smallsetminus F\subseteq U_{t}^{\prime\prime}\smallsetminus W_{t}^{\prime\prime}$. Hence $|E\smallsetminus F|\leqslant 1$, which by Lemma 4.2 contradicts that $(E,F)$ is a tetra-separation. ∎
@@
(i) $\Rightarrow$ (ii): this is 8.27.
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Assume that $G_{t}$ is bad. Then $(U_{t},W_{t})$ fails to be a tetra-separation, unless $G_{t}$ and both neighbouring bags are $K_{2}$’s, by 8.27. In this case, both vertices of $G_{t}$ lie in $U_{t}\smallsetminus W_{t}$, so $(U_{t},W_{t})$ is distinct from all tetra-separations in the tetra-star $\sigma(\mathcal{O})$. Therefore, $(U_{t},W_{t})$ interlaces $\sigma(\mathcal{O})$, and so $(U_{t},W_{t})$ is crossed by another tetra-separation, by Lemma 8.34. ∎
@@
We combine Lemmas 2.6.3 and 2.6.4 from [Tridecomp]. ∎
@@
By the (Crossing Lemma). (4.1), the centre $Z$ equals the vertex-centre and all links have size one. Let $\mathcal{O}$ be the Block-bagel induced by $(A,B)$ and $(C,D)$, which has centre $Z$. By Lemma 8.29, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$. By Lemma 8.34, the tetra-star $\sigma(\mathcal{O})$ is a splitting star of $N$. Since both $(A,B)$ and $(C,D)$ interlace both $\sigma$ and $\sigma(\mathcal{O})$, we obtain $\sigma=\sigma(\mathcal{O})$. Hence it remains to compute the torso $\tau$ of $\sigma(\mathcal{O})$. The torso $\tau$ is 4-connected by Lemma 8.36. We can obtain $\tau$ from $\mathcal{O}$ by replacing each good bag $G_{t}$ by a $K_{4}$ on the vertex set $Z\cup\{u,v\}$ where $u$ and $v$ are the adhesion-vertices of $G_{t}$, leaving only $K_{2}$-bags and triangle-bags, and contracting some edges in $K_{2}$-bags. Hence $\tau$ is a generalised wheel with centre $Z$. ∎

### v2

@@ 8.3. Totally-nested tetra-separations from good bags
Case $\pi=\{x,y\}$. We may assume that both $z_{1},z_{2}$ are vertices, by (6.4.1). Since the bag $G_{t(0)}$ is 2-connected, it contains two independent $x^{\prime}$$y^{\prime}$ paths. By Lemma˜8.21, these paths contain two independent $\hat{x}$$\hat{y}$ paths which are also included in the bag $G_{t(0)}$. We find a third $\hat{x}$$\hat{y}$ path $G$ that avoids both the bag $G_{t(0)}$ and the centre, by the (One-Path Lemma). (8.10). So we are done by (6.4.3).
Case $\pi=\{z_{1},z_{2}\}$. We find three independent $z^{\prime}_{1}$$z^{\prime}_{2}$ paths in GxyG-x-y by Lemma˜8.19. These contain three independent $\hat{z}_{1}$$\hat{z}_{2}$ paths by Lemma˜8.21. So we are done by (6.4.3).
Case $\pi=\{x,z_{1}\}$. We find an $x^{\prime}$$z^{\prime}_{1}$ path with interior in $\mathring{G}_{t(0)}$, by the (Two-Fan Lemma). (8.11). This path contains a $\hat{x}$$\hat{z}_{1}$ path $P_{1}$, by Lemma˜8.21. Next, we find two independent $x^{W}$$z^{\prime}_{1}$ paths $P_{2}$ and $P_{3}$ in $G$ that internally avoid $G_{t(0)}$, by Lemma˜8.20. Both paths $P_{2}$ and $P_{3}$ are $\hat{x}$$\hat{z}_{1}$ paths by Lemma˜8.16. Hence the $P_{i}$ witness (6.4.3).
@@
We claim that $z^{\prime}_{1}=z_{1}$. On the one hand, $z^{\prime}_{1}$ has $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma˜8.16, so $z^{\prime}_{1}$ remains in the separator when taking the right-reduction $(U^{\prime\prime},W^{\prime\prime})$ of $(U^{\prime},W^{\prime})$. However, $y^{\prime}$ becomes the edge $y$ of the $K_{2}$-bag $G_{t(1)}$ in the separator $S(U^{\prime\prime},W^{\prime\prime})$ by ˜8.15. Hence $z^{\prime}_{1}$ has two neighbours $y^{\prime}$ and $v$ in $U^{\prime\prime}\smallsetminus W^{\prime\prime}$. Thus $z^{\prime}_{1}$ also remains in the separator when taking the left-reduction $(U,W)$ of $(U^{\prime\prime},W^{\prime\prime})$. In particular, $z^{\prime}_{1}=z_{1}$.
@@
Suppose otherwise that $G_{t(0)}$ is 2-connected and has at least four vertices. We will find five independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G$ for (6.7.2). There exist two disjoint $\{z^{\prime}_{1},z^{\prime}_{2}\}$$\mathring{G}_{t(0)}$ paths $P_{1},P_{2}$ by Lemma˜8.8. Since $G_{t(0)}$ is 2-connected, we can extend the paths $P_{1},P_{2}$ to two disjoint $\{z^{\prime}_{1},z^{\prime}_{2}\}$$\{x^{\prime},y^{\prime}\}$ paths with interior in $\mathring{G}_{t(0)}$. We find three independent $\{x_{1}^{W},x_{2}^{W}\}$$\{z_{1}^{\prime},z_{2}^{\prime}\}$ paths in $G[W]$ by Lemma˜8.20. The five paths are independent, and they contain five $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths by Lemma˜8.21.
Case $\pi_{1}=\{x,z_{1}\}$ and $\pi_{2}=\{y,z_{2}\}$. Here it suffices to find five independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G$ for (6.7.2). There are two independent $x^{\prime}$$y^{\prime}$ paths in $G_{t(0)}$ by 2-connectivity. We find three independent $z^{\prime}_{1}$$z^{\prime}_{2}$ paths in $G[W]-\{x^{\prime},y^{\prime}\}$ by Lemma˜8.19. The latter three paths are internally disjoint from the former two. By Lemma˜8.21, the five paths contain five independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths.
@@
Assume $\pi=\{x,z_{1}\}$. The vertex $z^{\prime}_{1}$ in the centre has $\geqslant 3$ neighbours in $W^{\prime}\smallsetminus U^{\prime}$ by Lemma˜8.16. Hence $z^{W}_{1}=z^{\prime}_{1}$. Since the $\geqslant 3$ neighbours also lie in $W\smallsetminus U$ by ˜5.9, we further get that $z_{1}$ must be the edge in $\pi$. Since $z_{1}$ is an edge, $x$ is a vertex which has exactly one neighbour in $G[W]-z_{1}^{W}$. Hence $G_{t(-1)}$ is a $K_{2}$. But then $x$ is the edge of $G_{t(-1)}$ by ˜8.15, a contradiction.
Assume $\pi=\{x,y\}$. Say $x$ is a vertex and $y$ is an edge. Then $x=x^{\prime}$, so $G_{t(-1)}$ is 2-connected by ˜8.15. Hence $x$ has two neighbours $u,v$ in $G_{t(-1)}$. By the (Helpful Lemma). (8.6), neither $u$ nor $v$ is an endvertex of the edge $y$. Hence $\{x,y\}$ is not $W$-weird, a contradiction.
@@
Let $\{\pi_{1},\pi_{2}\}$ be a balanced bipartition of $S(U,W)$. Neither $\pi_{i}$ is $W$-weird by Lemma˜8.25. In particular, $\{\pi_{1},\pi_{2}\}$ is 2-potter-linked.
To show that $\{\pi_{1},\pi_{2}\}$ is 1-potter-linked, it suffices to find three independent $\hat{\pi}_{1}$$\hat{\pi}_{2}$ paths in $G[W]$. If one of the $\pi_{i}$ is of the form $\{x,y\}$, then we find three such paths by Lemma˜8.20 and Lemma˜8.21. Otherwise $z_{1}\in\pi_{1}$ and $z_{2}\in\pi_{2}$, say. Then we find three desired paths by Lemma˜8.19 and Lemma˜8.21. ∎
@@
We may assume the (Block-bagel bag setting). (8.14) with $t=t(0)$. Then $(U_{t},W_{t})$ is a tetra-separation by Lemma˜8.17. It suffices to show that $(U_{t},W_{t})$ is externally 5-connected, by ˜6.16. This follows from Lemma˜8.22 (half-connected), Lemma˜8.23 (3-linked), Lemma˜8.24 (0-potter-linked) and Lemma˜8.26 (1-potter-linked and 2-potter-linked). ∎
@@ 8.4. Splitting stars and torsos
The following two lemmata are supported by Figure˜8.7.
@@
Assume the (Block-bagel setting). (8.4). Let $P$ and $Q$ be the two paths that are the components of Oo(C)o(D)O-o(C)-o(D). Assume $o(A)\in P$ and $o(B)\in Q$. Let $c$ denote the end of $P$ that is adjacent or incident to $o(C)$ in $O$. Define $d$ similarly. Let $s$ be a node in the same component of Po(A)P-o(A) as $c$ such that $G_{s}=K_{2}$. Let $P^{\prime}:=dP\mathring{s}$ and $Q^{\prime}:=\mathring{s}Pc\cup Q$.
@@
Matching-condition: We have to show that, if $d$ in an edge, then it does not share an endvertex with $e$. The endvertices of $d$ lie in the corners for $AD$ and $BD$ since there do not exist dangling edges in the crossing-diagram of $(A,B)$ and $(C,D)$ by the (Crossing Lemma). (4.1). Both endvertices of $e$ lie in the corner for $AC$ by Lemma˜8.2 (i). Hence $d$ and $e$ share no ends.
@@
If $d$ is a vertex, then we claim that every neighbour $u$ of $d$ in $A\smallsetminus B$ also also lies in $A^{\prime}\smallsetminus B^{\prime}$. Indeed, since $ud$ is not a dangling edge, $u$ is contained in the corner $AD$ or the $A$-link with respect to $(A,B),(C,D)$. By Lemma˜8.2 (i), no endvertex of $e$ lies in the $A$-link with respect to $(A,B),(C,D)$. So, we have $u\in A^{\prime}\smallsetminus B^{\prime}$.
It remains to show that both vertices $z_{1}^{\prime}$ and $z_{2}^{\prime}$ in the centre have at least two neighbours in $A^{\prime}\smallsetminus B^{\prime}$. Let $s^{\prime}$ be the neighbour of $s$ on $P^{\prime}$. Observe that either $s^{\prime}$ or $ss^{\prime}$ is helpful and belongs to the corner $AC$. Moreover, by the (Helpful Lemma). (8.6) there is a helpful element that belongs to the $AD$-corner. Hence, there are at least two helpful elements that belong to $(ss^{\prime})\cup T_{AD}$, where $T_{AD}$ is the $AD$-block-path. By Lemmas˜8.7 and 8.8, we are done.
@@
Assume the (Block-bagel setting). (8.4). Let $P$ and $Q$ be the two paths that are the components of Oo(C)o(D)O-o(C)-o(D). Assume $o(A)\in P$ and $o(B)\in Q$. Let $c$ denote the end of $P$ that is adjacent or incident to $o(C)$ in $O$. Define $d$ similarly. Let $e$ be an edge of $O$ with 2-connected bags at both ends. Assume that $e$ stems from the $AC$-block-path. Let $P^{\prime}:=dP\mathring{e}$ and $Q^{\prime}:=\mathring{e}Pc\cup Q$. Then:
@@
If the $A$-link or the $C$-link contains a vertex, then this vertex is contained in a 2-connected bag $G_{t}$ with $t\in V(T^{\prime})$, by Lemma˜8.2. Otherwise both the $A$-link and the $C$-link consist of edges. Then $G_{o(A)}$ and $G_{o(C)}$ are $K_{2}$-bags. Each of the two bags $G_{o(A)},G_{o(C)}$ has exactly one endvertex $v_{A},v_{C}$ in the corner for $AC$ by Lemma˜8.2, respectively. If $T^{\prime}$ is empty, then $v_{A}=v_{C}$, and $A\cap C=\{v_{A}\}\cup Z^{\prime}$ has size three. ∎
@@
Note that the existence of $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$ follows from $(A,B)$ and $(C,D)$ exhibiting properties (i)(iii). Let $T$ denote the $A^{\prime}C^{\prime}$-block-path, and let $T^{\prime}:=T-o(A^{\prime})-o(C^{\prime})$. By Lemma˜8.32, $T^{\prime}$ is non-empty. Let $\mathcal{O}^{\prime}=\mathcal{O}$ be the block-bagel induced by $(A^{\prime},B^{\prime})$ and $(C^{\prime},D^{\prime})$.
@@
    $S(A^{\prime\prime},B^{\prime\prime})=\{e\}\cup Z^{\prime}\cup\text{($D^{\prime}$-link)}$, and
@@
    $S(C^{\prime\prime},D^{\prime\prime})=\{e\}\cup Z^{\prime}\cup\text{($B^{\prime}$-link)}$, and
@@
Assume for contradiction that $\sigma(\mathcal{O})$ is interlaced by a totally-nested tetra-separation $(E,F)$ of $G$. By Lemma˜8.33, there is a 2-connected bag $G_{t}$ of $\mathcal{O}$ with $E\subseteq U_{t}^{\prime}$, say. By Lemma˜5.7, we have $(E,F)\leqslant(U_{t}^{\prime\prime},W_{t}^{\prime\prime})$ where $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$ is the right-reduction of $(U^{\prime}_{t},W^{\prime}_{t})$.
Assume first that $G_{t}$ is good. Then $(U_{t},W_{t})\in\sigma(\mathcal{O})$ by definition of $\sigma(\mathcal{O})$. Since $(E,F)$ interlaces $\sigma(\mathcal{O})$, we have $(U_{t},W_{t})<(E,F)$ or $(U_{t},W_{t})<(F,E)$. We cannot have $(U_{t},W_{t})<(F,E)$ since otherwise $\emptyset\neq E\smallsetminus F\subseteq(W_{t}\smallsetminus U_{t})\cap(U_{t}^{\prime\prime}\smallsetminus W_{t}^{\prime\prime})$ contradicts that $U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}\subseteq U_{t}\smallsetminus W_{t}$ by ˜5.2. So $(U_{t},W_{t})<(E,F)\leqslant(U_{t}^{\prime\prime},W_{t}^{\prime\prime})$. Since $(U_{t},W_{t})$ is the left-reduction of $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$, some vertex in $S(E,F)$ must have $\leqslant 1$ neighbour in $E\smallsetminus F$, contradicting the degree-condition.
Otherwise, $G_{t}$ is bad. Since $G_{t}$ is 2-connected, it is a triangle, and since $G_{t}$ is bad, both neighbouring bags in $\mathcal{O}$ are 2-connected. Hence both adhesion-vertices of $G_{t}$ are contained in the separator of $(U^{\prime\prime}_{t},W^{\prime\prime}_{t})$. Therefore, the unique interior vertex of the triangle-bag $G_{t}$ is the only vertex of $U^{\prime\prime}_{t}\smallsetminus W^{\prime\prime}_{t}$. Since $(E,F)\leqslant(U_{t}^{\prime\prime},W_{t}^{\prime\prime})$, we have $E\smallsetminus F\subseteq U_{t}^{\prime\prime}\smallsetminus W_{t}^{\prime\prime}$. Hence $|E\smallsetminus F|\leqslant 1$, which by Lemma˜4.2 contradicts that $(E,F)$ is a tetra-separation. ∎
@@
˜(i) $\Rightarrow$ ˜(ii): this is ˜8.27.
$\neg$ ˜(i) $\Rightarrow$ $\neg$ ˜(ii). Assume that $G_{t}$ is bad. Then $(U_{t},W_{t})$ fails to be a tetra-separation, unless $G_{t}$ and both neighbouring bags are $K_{2}$’s, by ˜8.27. In this case, both vertices of $G_{t}$ lie in $U_{t}\smallsetminus W_{t}$, so $(U_{t},W_{t})$ is distinct from all tetra-separations in the tetra-star $\sigma(\mathcal{O})$. Therefore, $(U_{t},W_{t})$ interlaces $\sigma(\mathcal{O})$, and so $(U_{t},W_{t})$ is crossed by another tetra-separation, by Lemma˜8.34. ∎
@@
We combine Lemmas 2.6.3 and 2.6.4 from [CK23]. ∎
@@
By the (Crossing Lemma). (4.1), the centre $Z$ equals the vertex-centre and all links have size one. Let $\mathcal{O}$ be the Block-bagel induced by $(A,B)$ and $(C,D)$, which has centre $Z$. By Lemma˜8.29, the tetra-star $\sigma(\mathcal{O})$ is a star of totally-nested tetra-separations of $G$. By Lemma˜8.34, the tetra-star $\sigma(\mathcal{O})$ is a splitting star of $N$. Since both $(A,B)$ and $(C,D)$ interlace both $\sigma$ and $\sigma(\mathcal{O})$, we obtain $\sigma=\sigma(\mathcal{O})$. Hence it remains to compute the torso $\tau$ of $\sigma(\mathcal{O})$. The torso $\tau$ is 4-connected by Lemma˜8.36. We can obtain $\tau$ from $\mathcal{O}$ by replacing each good bag $G_{t}$ by a $K_{4}$ on the vertex set $Z\cup\{u,v\}$ where $u$ and $v$ are the adhesion-vertices of $G_{t}$, leaving only $K_{2}$-bags and triangle-bags, and contracting some edges in $K_{2}$-bags. Hence $\tau$ is a generalised wheel with centre $Z$. ∎